The binding energy of electron in a metal is 193kj/mol. find the threshold frequency of the metal?

Grade:11

1 Answers

somnath mondal
6 years ago
From the photoelectric effect of Hertz....E=K.E+P.E....this potential energy is the binding energy...again P.E= Work functionP.E=h*threshold frequencyThus; threshold frequency =P.E/h =(193 J/mol)/(6.626*10^-34 J.s). =(2.91*10^38 s^-1.mol^-1)/(6.023*10^23 mol^-1). =4.8*10^14 s^-1